OpenStudy (anonymous):

Two parallel conducting plates are connected to a constant voltage source. The magni- tude of the electric field between the plates is 2429 N/C. If the voltage is tripled and the distance between the plates is reduced to 1/5 the original distance, what is the magnitude of the new electric field?

OpenStudy (anonymous):

It is E=U/d. So the new electric field is 3U/(d/5)=15U/d=15E=15*2429N/C=25504.5N/C